OpenStudy (anonymous):

Find the perimeter of the rectangle in simplest from. Width= 2√6 and Length= 4+ √6.. Please help

OpenStudy (lgbasallote):

perimether = length + length + width + width \[P = (4 + \sqrt 6) + (4 + \sqrt 6) + 2\sqrt 6 + 2\sqrt 6\] \[P = 4 + \sqrt 6 + 4 + \sqrt 6 + 2\sqrt 6 + 2\sqrt 6\] \[P = 4 + 4 + \sqrt 6 + \sqrt 6 + 2\sqrt 6 + 2\sqrt 6\] do you see it now?

OpenStudy (anonymous):

yes... thanks :)

OpenStudy (lgbasallote):

wonderful!